# JEL: C81, O32

### v1

@@ 3.2. Objective
3.1 embeds several restrictions, intended to yield a measure of new investment in information relevant to human-facing medical care. We view this set of studies as the set most likely to generate information relevant to human mortality. Studies involving animals, literature summaries, and re-analyses of existing data, for example, are excluded. We take the “potential universe” plotted in Figure 1 as our baseline sample of potential clinical trials. We further restrict attention to the 1,821,429 records published in or after 2010.161616There is a trade-off between data coverage and quality. Around 2010, a set of policy changes—including trial registration requirements and harmonization in publication formatting—affected trial disclosure in scientific publications. See Laine et al., (2007) and Schulz et al., (2010) for examples of two such changes.
@@
To quantify our ability to identify the clinical trials in this set of publications, we hand-label approximately 3,000 randomly selected publications based on the content of their abstracts. [^17] We develop a custom labelling interface that reduces the time required to label 100 records, for the authors, by a factor of seven. Section B.1 describes this tool and its application to this context in detail. Of the hand-labelled publications, 11.2% meet the criteria for inclusion in our sample. The hand-labelled data are split into three subsets—validation, training, and testing—based on their eventual use.
@@
As a baseline, we assess the performance of several standard machine learning algorithms. The results are displayed in Figure 2. Further details on the construction of this figure are given in Section B.2. Each algorithm is estimated on the hand-labels assigned to the publications in the training and validation sample splits. For feature vectors, we use either TF-IDF embeddings computed in the corpus of abstracts in the hand-labeled sample or the embeddings of each abstract obtained from the SentenceTransformer language model (Reimers and Gurevych,, 2019). We find that the performance of the standard machine learning algorithms is again unsuitable for our application. At a 90% true positive rate, the best performing model identifies 50 true positives for every 50 false positives.
Figure 2. Performance of Standard Machine Learning Methods
@@
Notes: Figure 2 displays the receiver operating characteristic of several machine learning models, trained on two types of embeddings, for classifying whether a publication satisfies the restrictions enumerated in 3.1. The hyper-parameters of each model are determined with 10-fold cross-validation. The models are trained in the training and validation data set. Error rates are estimated in the testing data set. In addition, we display the performance of our, ensemble, fine-tuned large language model in purple. This is the same curve labelled “Ensemble” in Figure 3.
@@ 3.3. Prompt Design and Fine-Tuning
We construct a large language model optimized for our task. This is accomplished in three steps. First, we iteratively construct a set of prompts that exhibit good performance when posed to proprietary models—OpenAI’s GPT-3.5 and GPT-4. Second, we extract noisy labels for a large number of publications in our sample by querying the proprietary models. Third, the noisy labels are used to train an off-the-shelf large language model. The resulting model is then used to identify clinical trials in our baseline sample. This process of constructing specialized large language models is referred to as “fine-tuning” (see e.g., Taori et al.,, 2023, for a prominent application of this idea).
@@ 3.3.1. Prompt Design
We devise three types of prompts and measure their performance on samples of abstracts taken from our validation dataset. We measure model performance with the hand-labeled data, conduct error analyses, and iteratively revise prompts. When modifications to the prompt cease to yield improvements in performance, we finalize prompts best suited for GPT-3.5 and GPT-4, respectively. Section B.3 provides more detail on this process. Figure 3 displays estimates of the true and false positive rates computed in the test data, for both proprietary models, with black dots.
Figure 3. Receiver Operating Characteristic
@@
Notes: Figure 3 displays estimates of the receiver operating characteristic of several models used to classify the publications indexed by PubMed / MEDLINE according to whether they satisfy the restrictions enumerated in 3.1. All metrics are computed in the testing split of the hand-labelled publications. Estimates of the true positive rate and false positive rate of OpenAI’s proprietary models GPT-3 and GPT-4 are indicated with black dots. The curves give the performance of four fine-tuned, open-source large language models, in addition to an ensemble model. The blue dots indicate the true positive rate and false positive rate of the models used to construct the “Conservative,” “Moderate,” and “Liberal” samples of clinical trials.
@@ 3.3.2. Fine-Tuning
We compute noisy labels for 64,000 randomly selected publications in our baseline sample using the best performing prompts for both GPT-3 and GPT-4. These noisy labels are used as data to train off-the-shelf language models from two architecture classes: (1) RoBERTa (Liu et al.,, 2019) and (2) BioMedBERT (Gu et al.,, 2021). Both models are available in either “baseline” or “large” varieties. [^19] There are many open-source large language models (e.g., LLaMa, Mistral, Pythia). Many are several orders of magnitude (e.g., 7-70 billion parameters) larger than RoBERTA ($\sim$ 350 million parameters). We selected RoBERTA and BioMedBERT via trial-and-error that included testing the performance of these larger models. In our setting, RoBERTA outperforms LLaMA (both 7 and 70 billion parameter versions, trained with QLoRA). Both models exhibited comparable performance, but RoBERTa was substantially faster and simpler to train. The BioMedBERT model is an instance of a BERT model (Devlin et al.,, 2019), that has itself been fine-tuned on the abstracts of the publications indexed in PubMed. We systematically vary model architecture, size, and pre-training regimen. We observe no large differences in performance stemming from these hyper-parameters. Sections B.4 and B.5 give further details the fine-tuning process and the results of these experiments, respectively.
@@ 3.4. Performance
Given the text of an abstract, our fine-tuned language models output a probability that the publication satisfies the restrictions enumerated in 3.1. Publications whose probabilities fall above a chosen threshold are classified as belonging in our sample. Figure 3 displays estimates of the true and false positive rates, computed with the test data, as we vary this threshold for both sizes of the RoBERTa and BioMedBERT models. [^20] Each of the models whose results are displayed in Figure 3 are trained with noisy labels obtained with GPT-4. We document in Section B.5 that models trained with noisy labels obtained from GPT-3.5 or with the 1000 hand-labeled records in the training data perform significantly worse. The figure additionally displays the performance of an ensemble model estimated in the training data. This model is obtained by estimating a logistic regression of the hand-labels on the probabilities output by all four models displayed Figure 3. The ensemble model is used to produce our final sample.
@@
The fine-tuned models are able to match the performance of GPT-4 in the test data. We choose three thresholds according to the stringency with which they enforce the sample restrictions. The estimated true and false positive rates associated with these points are displayed in Figure 3 and labeled as “Conservative,” “Moderate,” and “Liberal.” Our preferred sample is associated with the conservative threshold. [^21] The liberal model may be particularly useful for conducting literature reviews, where a near-perfect true positive rate is needed. For every 82 true positives, the conservative model identifies 18 false positives. [^22] In the test data, the conservative model assigns incorrect labels to 27 of 993 papers. We conduct an error analysis. See Section B.6 for further details. In 13 cases, there is a clear error. In 14 cases, however, errors are associated with records that were difficult to categorize for a human labeller. For example, PubMed record 32737793 is flagged as satisfying 3.1 with all three thresholds, but is a literature review. By contrast, PubMed record 27880726 was categorized as satisfying 3.1 twice by a human labeller. It is excluded from both the conservative and moderate model-generated samples. On inspection, the abstract does not explicitly state that the study enrolled human subjects, but hints that it may have been conducted in an animal model. Review of the associated full text confirms that this study did, in fact, enroll only rats. We report results associated with the moderate and liberal thresholds in the appendix as tests of robustness. Our final, conservative, sample consists of 152,027 publications classified as satisfying the restrictions enumerated in 3.1.
@@
Appendix A compares the contents of our sample to the counts of potential clinical trials plotted in Figure 1. Although certain elements of PubMed metadata—including the union of all NLM tags—capture many of the records in our final sample, we confirm that they miss many records that we flag as trials and include many records that do not satisfy 3.1. No combination of existing search strategies, then, classifies records with the same accuracy or precision as our final sample.
4. Medical Research as a Productivity Indicator
Figure 1 suggests that, across metrics, there has been a substantial increase in the quantity of medical research over time, including since 2010. This is consistent with documented increases in the quantity of scientific papers in Bloom et al., (2020) and Park et al., (2023), as well as economy-wide increases in research effort reported in Goldin et al., (2024).
4.1. Decomposing Medical Research
We decompose this pattern, into changes attributable to high-productivity, high-capital investments—clinical trials—and changes attributable to other forms of clinical research. Our census of clinical trials provides measures for the first category. Although there are many reasonable ways to construct samples of other clinical research, our preferred approach defines this category as those scientific publications that cite clinical trials. [^23] Alternative approaches to construction of this sample, as in Figure 1, also suggest an increase in the quantity of non-clinical trial research. We prefer to use this citation-based definition because it allows for consistency in sample definition. Other approaches that rely on keyword searches, NLM tags, etc., as Section 3 and its appendices make clear, capture different sets of records over time. These are records of research investment that necessarily engage with findings from trials (such as meta-analyses and re-analyses of existing data), but, as discussed in Section 2, are essentially always lower productivity and less capital-intensive.
@@ 4.1. Decomposing Medical Research
Figure 4 documents that the quantity of clinical trials—the black line—has remained constant over the past ten years. In contrast, other clinical research has increased substantially. Observe that we display the percent growth in the quantity of clinical trials in our sample published in each calendar year, alongside the growth in the set of papers that cite clinical trials. Each series is normalized to begin at zero. To account for concerns about citation delays and truncation, we construct various measures of papers that cite clinical trials, using 2–6 year citation counts. Two-year citation counts, thus, are available starting in the year 2012 (two years into our sample) and count publications that cite clinical trials published in the preceding two years. [^24] Figure C.3 and Figure C.4, displayed in the supplemental appendix, document that this trend is robust to alternative sample definitions, using either our liberal or moderate samples, and to citation-weighting count data. This presentation masks substantial differences in the levels of these series. There are 10,903 clinical trials in our census in 2010, and 73,253 publications that cite clinical trials (using two-year citations) in 2012.
Figure 4. Growth in Clinical Research, Stability of Clinical Trials
@@
Notes: Figure 4 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year. Each series is reported in terms of the percent change relative to its initial value. The sample of published clinical trials is constructed with the conservative model. To address truncation, we report the number of publications that cite clinical trials published in the preceding $t$ years for each $t$ between 2 and 6.
4.2. Stable Composition of Clinical Trials
To contextualize the trends displayed in Figure 4, it is helpful to return to the simple model from Section 2. We document that the quantity of capital-intensive, high-productivity research inputs remains unchanged over our period of interest, while low-capital, low-productivity inputs increase. As changes in outputs in this sector (mortality) are assumed to remain constant (see Section 2), there are thus three ways of reconciling these facts in our stylized framework: (i) other clinical research is sufficiently low capital or low productivity, across the board, as to have essentially no impact on growth, (ii) increases in other clinical research have, heterogeneously, low capital or low productivity, or (iii) clinical trial research productivity is declining over this time period, fully offsetting the growth in other types of research. If either Case (i) or Case (ii) were an accurate characterization, rising observed research effort—increases in the number of scientific publications or the number of records in PubMed that the previous methods have identified as clinical trials—need not imply diminishing returns to high-capital investment.
Our data allow us to reject the hypothesis that we are in Case (iii). Figure 5 documents that—on four dimensions—measures of trial productivity, across clinical trials, have remained stable over the past decade. Panel A of Figure 5 considers the share of publications in our census with the following characteristics: any funding from a government agency (purple), any (three-year) citations from a leading journal in medicine (teal), or any citations from a leading journal conditional on having public funding (green). [^25] We designate a publication as having public funding if it appears in the National Institutes of Health RePORTER data, as linked to a funded research grant, or if PubMed reports a source of research funding, including funding from non-U.S. agencies. To designate a set of the “top” journals in medicine, we follow Angrist et al., (2020) We collect all citations originating from the “trunk journals” in medicine, the Journal of the American Medical Association and the New England Journal of Medicine, to records in PubMed between 2010 and 2022. We designate a journal as being “leading” of it received at least 100 citations from a trunk journal over this time period. This yields a list of 84 journals. See Section A.3 for details. Appendix C reproduces this figure using five-year citations. The trend is unchanged. The trends in Panel A communicate two findings. First, all three proxies for research productivity are stable, essentially unchanging, over our time period. Second, nearly 55 percent of clinical trials are never cited by a leading journal, suggesting substantial, but time-invariant, heterogeneity across trials.
Figure 5. Stability of Heterogeneity
Panel A: Public Funding and Low-Quality Research
Panel B: The Composition of Quality over Time

### v2

@@ 3.2. Objective
3.1 embeds several restrictions, intended to yield a measure of new investment in information relevant to human-facing medical care. We view this set of studies as the set most likely to generate information relevant to human mortality. Studies involving animals, literature summaries, and re-analyses of existing data, for example, are excluded. We take the “potential universe” plotted in Figure 3.1 as our baseline sample of potential clinical trials.
We further restrict attention to the 1,821,429 records published in or after 2010, as there is a trade-off between data coverage and quality in this setting. In the 1990s and 2000s, medical experts created a standardized system for clinical trial reporting in scientific publications. By the time of the most recent Consolidated Standards of Reporting Trials (CONSORT) in 2010, nearly all medical journals required authors to comply with a standardized structure for all clinical trial publications (Schulz et al.,, 2010). As a result, most trial publications have a distinctive format and include a pre-defined set of elements. Although our procedure can be applied to any type of scientific publication text, the standardization required by the CONSORT Statement changes the relationship between clinical trial research and publications in the early 2000s. To answer the research question that motivates this paper—is the quantity of medical research rising over time?—we elect to proceed with a narrower sample to avoid this issue of interpretability.
@@
To quantify our ability to identify the clinical trials in this set of publications, we hand-label approximately 3,000 randomly selected publications based on the content of their abstracts. We develop a custom labeling interface that reduces the time required to label 100 records, for the authors, by a factor of seven. Section C.1 describes this tool and its application to this context in detail. Of the hand-labeled publications, 11.2% meet the criteria for inclusion in our sample. The hand-labeled data are split into three subsets—validation, training, and testing—based on their eventual use.
@@
As a baseline, we assess the performance of several standard machine learning algorithms. The results are displayed in Figure 3.2. Further details on the construction of this figure are given in Section C.2. Each algorithm is estimated on the hand-labels assigned to the publications in the training and validation sample splits. For feature vectors, we use either TF-IDF embeddings computed in the corpus of abstracts in the hand-labeled sample or the embeddings of each abstract obtained from the SentenceTransformer language model (Reimers and Gurevych,, 2019). We find that the performance of the standard machine learning algorithms is again unsuitable for our application. At a 90% true positive rate, the best performing model identifies 50 true positives for every 50 false positives.
Figure 3.2. Performance of Standard Machine Learning Methods
@@
Notes: Figure 3.2 displays the receiver operating characteristic of several machine learning models, trained on two types of embeddings, for classifying whether a publication satisfies the restrictions enumerated in 3.1. The hyper-parameters of each model are determined with 10-fold cross-validation. The models are trained in the training and validation data set. Error rates are estimated in the testing data set. In addition, we display the performance of our, ensemble, fine-tuned large language model in purple. This is the same curve labelled “Ensemble” in Figure 3.3.
@@ 3.3. Prompt Design and Fine-Tuning
We construct a large language model optimized for our task. This is accomplished in three steps. First, we iteratively construct a set of prompts that exhibit good performance when posed to proprietary models—OpenAI’s GPT-3.5 and GPT-4. A priori, we expect GPT-4 to output labels of slightly lower quality than hand-labeled output and GPT-3.5 to produce slightly noisier labels than GPT-4. The extraction of GPT-3.5 labels is substantially faster and cheaper. Second, we extract noisy labels for a large number of publications in our sample by querying the proprietary models. Third, the noisy labels are used to train an off-the-shelf large language model. The resulting model is then used to identify clinical trials in our baseline sample. This process of constructing specialized large language models is referred to as “fine-tuning” (see e.g., Taori et al.,, 2023, for a prominent application of this idea).
@@ 3.3.1. Prompt Design
To produce these labels at scale, each of the two models must be appropriately prompted. That is, each pre-trained model must be provided with a block of text as an input and asked to return the appropriate completion of this input. A priori, it is not clear what prompt structure will work well. The relative infancy of this area of research renders it difficult to identify a set of “best practices.” [^34] See, for one example of this evolving area of work, a guide to “prompt engineering” from OpenAI: https://platform.openai.com/docs/guides/prompt-engineering. All of the prompts that we consider are displayed as figures in Appendix E.
We identify three general prompt formats, which differ both in the amount of detail provided about our classification task and in the structure of the requested model completion. The simplest prompt provides a version of our sample definition, 3.1, and the text of an abstract, and asks the model to return ‘TRUE’ if the abstract satisfies these criteria. Otherwise, it returns ‘FALSE’. The text of this prompt is displayed in Figure E.1. We refer to this prompt as Prompt 1.0. A second, more complicated prompt provides the same definition, with a set of examples of publication characteristics that do and do not satisfy the definition. Here, the prompt asks the model to return either ‘TRUE’ or the name of a specific excluded category. The text of this prompt is displayed in Figure E.7. We refer to this prompt as Prompt 2.0. A third, more complex prompt provides the same definitions and examples, but asks the model to return either ‘TRUE’ or an explanation of why the record does or does not satisfy our sample definition. The text of this prompt is displayed in Figure E.10. We refer to this prompt as Prompt 3.0. We devise initial language for each prompt iteratively, using small numbers of records from our hand-labelled validation data.
We test each of these three prompts in our 1000-record hand-labelled validation dataset, using both GPT-3.5 and GPT-4. We conduct a detailed error analysis, reported in LABEL:tab:Error-Analysis. For each instance in which a model returns a label (TRUE/FALSE) that differs from that in the hand-labelled dataset, we inspect the record. We categorize errors into types and sub-types. Details are given in Section C.3.
This exercise is instructive on three margins. First, it highlights the highest performing prompts. Second, it indicates particular types of errors in categorization—which suggest opportunities for more precise language in a prompt. Third, it draws attention to differences in the performance of the two models. Observe that, in the “Other” error type, we include a sub-type called “overly literal interpretation of inclusion criteria.” We primary record errors of this type for Prompt 3, which asked the model to return a completion that described why, or why not, a record was classified as being in our sample. Here, GPT-3.5 makes 30 such errors—classifying records as FALSE by recapitulating the sample definition provided. GPT-4 makes two errors. The insight that emerges from this exercise, then, is that different prompt structures may be preferable depending on the model used, and that, by and large, GPT-4 is better suited for tasks that require (the resemblance of) more sophisticated reasoning.
We revise each class of prompt based on these findings. For Prompt 1.0, the simplest true/false prompt, we consider three variants. We refer to these revised prompts as Prompts 1.1, 1.2, and 1.3. The text of these prompts is displayed in Figures E.2, E.3, and E.5. For Prompts 2 and 3, we consider one variant each. We refer to these revised prompts as Prompts 2.1 and 3.1. The text of these prompts is displayed in Figures E.8 and E.11. These changes reflect the differences in performance catalogued in LABEL:tab:Error-Analysis.
Table 3.1 reports estimates of the true positive rate and false positive rate for each prompt, in both models, computed in the validation sample. Observe—for example, with Prompts 1.1, 1.2, and 1.3 queried to GPT-4—that even small changes in the text of a prompt yield substantial differences in performance. We attempt a second round of prompt iteration (Prompt 1.3). We observe that performance deteriorates with even small modification. Thus, we select the two highest performing prompts—one for each proprietary model. We use Prompt 2.0 to extract weak labels using GPT-3.5, and Prompt 1.2 to extract weak labels using GPT-4. Figure 3.3 displays estimates of the true and false positive rates computed in the test data, for both proprietary models, with black dots. These prompts are used to extract 64,000 labels—TRUE/FALSE flags—using GPT-3.5 and GPT-4. We refer to these as our set of “noisy labels.”
Table 3.1. Prompt Performance in Validation Data
Panel A: GPT-3.5
Prompt Type Prompt Sub-Type False Positive Rate True Positive Rate
1 0 0.247 0.788
1 0.172 0.898
2 0.037 0.584
3 0.037 0.489
2 0 0.162 0.876
1 0.176 0.905
3 0 0.248 0.722
1 0.171 0.883
Panel B: GPT-4
Prompt Type Prompt Sub-Type False Positive Rate True Positive Rate
1 0 0.202 0.971
1 0.065 0.949
2 0.049 0.934
3 0.056 0.912
2 0 0.081 0.964
1 0.072 0.964
3 0 0.167 0.971
1 0.068 0.956
Notes: Table 3.1 records the performance of each of eight prompt variants in a sample of 1,000 validation dataset records. Panel A reports performance associated with the proprietary model GPT-3.5. Panel B reports analogous statistics for model GPT-4. Prompt 1 asks the model to return TRUE or FALSE. Prompt 2 asks the model to return TRUE or the name of a specific, excluded category. Prompt 3 asks the model to return TRUE or an explanation of why the record should be excluded. Prompt sub-types correspond to various iterations. Sub-type 0 is the initial version of the prompt. Sub-types 1-3, where applicable, are subsequent iterations. Appendix E records the text of each prompt. We use Prompt 2, Sub-Type 0, to extract weak labels using GPT-3.5, and Prompt 1, Sub-Type 2 to extract weak labels using GPT-4. The performance measurements for these prompts are displayed in boldface.
Figure 3.3. Receiver Operating Characteristic
@@
Notes: Figure 3.3 displays estimates of the receiver operating characteristic of several models used to classify the publications indexed by PubMed / MEDLINE according to whether they satisfy the restrictions enumerated in 3.1. All metrics are computed in the testing split of the hand-labelled publications. Estimates of the true positive rate and false positive rate of OpenAI’s proprietary models GPT-3 and GPT-4 are indicated with black dots. The curves give the performance of four fine-tuned, open-source large language models, in addition to an ensemble model. The blue dots indicate the true positive rate and false positive rate of the models used to construct the “Conservative,” “Moderate,” and “Liberal” samples of clinical trials.
@@ 3.3.2. Fine-Tuning
We compute noisy labels for 64,000 randomly selected publications in our baseline sample using the best performing prompts for both GPT-3 and GPT-4. These noisy labels are used as data to train off-the-shelf language models from two architecture classes: (1) RoBERTa (Liu et al.,, 2019) and (2) BioMedBERT (Gu et al.,, 2021). Both models are available in either “baseline” or “large” varieties. [^36] There are many open-source large language models (e.g., LLaMa, Mistral, Pythia). Many are several orders of magnitude (e.g., 7-70 billion parameters) larger than RoBERTA ($\sim$ 350 million parameters). We selected RoBERTA and BioMedBERT via trial-and-error that included testing the performance of these larger models. In our setting, RoBERTA outperforms LLaMA (both 7 and 70 billion parameter versions, trained with QLoRA). Both models exhibited comparable performance, but RoBERTa was substantially faster and simpler to train. The BioMedBERT model is an instance of a BERT model (Devlin et al.,, 2019), that has itself been fine-tuned on the abstracts of the publications indexed in PubMed. We systematically vary model architecture, size, and pre-training regimen. We observe no large differences in performance stemming from these hyper-parameters. Sections C.4 and C.5 give further details concerning the fine-tuning process and the results of these experiments, respectively.
@@ 3.4. Performance
Given the text of an abstract, our fine-tuned language models output a probability that the publication satisfies the restrictions enumerated in 3.1. Publications whose probabilities fall above a chosen threshold are classified as belonging in our sample. Figure 3.3 displays estimates of the true and false positive rates, computed with the test data, as we vary this threshold for both sizes of the RoBERTa and BioMedBERT models. [^37] Each of the models whose results are displayed in Figure 3.3 are trained with noisy labels obtained with GPT-4. We document in Section C.5 that models trained with noisy labels obtained from GPT-3.5 or with the 1000 hand-labeled records in the training data perform significantly worse. The figure additionally displays the performance of an ensemble model estimated in the training data. This model is obtained by estimating a logistic regression of the hand-labels on the probabilities output by all four models displayed Figure 3.3. The ensemble model is used to produce our final sample.
@@
The fine-tuned models are able to match the performance of GPT-4 in the test data. We choose three thresholds according to the stringency with which they enforce the sample restrictions. The estimated true and false positive rates associated with these points are displayed in Figure 3.3 and labeled as “Conservative,” “Moderate,” and “Liberal.” Our preferred sample is associated with the conservative threshold. [^38] The liberal model may be particularly useful for conducting literature reviews, where a near-perfect true positive rate is needed. For every 82 true positives, the conservative model identifies 18 false positives. [^39] In the test data, the conservative model assigns incorrect labels to 27 of 993 papers. We conduct an error analysis. See Section C.6 for further details. In 13 cases, there is a clear error. In 14 cases, however, errors are associated with records that were difficult to categorize for a human labeller. For example, PubMed record 32737793 is flagged as satisfying 3.1 with all three thresholds, but is a literature review. By contrast, PubMed record 27880726 was categorized as satisfying 3.1 twice by a human labeller. It is excluded from both the conservative and moderate model-generated samples. On inspection, the abstract does not explicitly state that the study enrolled human subjects, but hints that it may have been conducted in an animal model. Review of the associated full text confirms that this study did, in fact, enroll only rats. We report results associated with the moderate and liberal thresholds in the appendix as tests of robustness. Our final, conservative, sample consists of 152,027 publications classified as satisfying the restrictions enumerated in 3.1.
@@
Each table displayed in Appendix A compares the contents of our sample to the counts of potential clinical trials plotted in Figure 3.1. Although certain elements of PubMed metadata—including the union of all NLM tags—capture many of the records in our final sample, we confirm that they miss many records that we flag as trials and include many records that do not satisfy 3.1. No combination of existing search strategies, then, classifies records with the same accuracy or precision as our final sample.
4. Disentangling Trends in Medical Research
Figure 3.1 indicates that, across measures, the total quantity of medical research appears to be increasing, including since 2010.404040We find similar patterns using alternative cuts of records indexed in PubMed. The total number of unique authors listed on scientific papers, the total number of citations given and received by scientific papers, the total number of scientific journals, and the total number of pages in scientific journals all exhibit similar patterns. These patterns are consistent with documented increases in the quantity of scientific papers in Bloom et al., (2020) and Park et al., (2023), as well as economy-wide increases in research effort reported in Goldin et al., (2024).
In this section, we use our census of clinical trials to decompose this trend. We contrast trends in the production of clinical trials with those documented for other forms of medical research. Differences between these series are informative in two ways. First, existing estimates of the quantity of medical research, such as those plotted in Figure 3.1, capture some combination of these two series (e.g., Bloom et al.,, 2020; Park et al.,, 2023; Chu and Evans,, 2021). Second, any proxy for the quantity of medical research is difficult to interpret if its relationship to the underlying research changes over time. This decomposition highlights the limitations of empirical strategies that treat counts of scientific publications as a proxy for research quantity.
There are many potential ways to construct samples of other (i.e., non-clinical trial) medical research. Our preferred approach defines this category as the collection of publications that cite clinical trials. (We confirm in Section 4.3.1 that this set includes only medical research.) Conceptually, of course, this “citing” sample is a subset of non-clinical trial medical research. Alternative methods of constructing a measure of “other” research introduce additional, complicating conceptual problems. Approaches that rely on keyword searches, NLM tags, etc., as Section 3 and its appendices make clear, capture different sets of records over time, as a consequence of changes in database structure. Approaches that examine other publications in the same journals as clinical trials will mechanically suggest shifts in total quantities, if the number of pages in journals change over time or if certain journals become online-only. [^41] Ioannidis et al., (2023) document that the number of medical journals is increasing. In 2024, the Annals of Emergency Medicine, the highest impact journal in the medical sub-field of emergency medicine, shifted to an online-only format. Brainard, (2020) discusses dozens of instances in which medical journals transitioned to online-only formats, then ceased publication entirely.
The primary challenge in using a citation-based measure is one of truncation: constructing measures of papers that cite clinical trials requires determining the window in which such citation must have occurred. To account for concerns about citation delays, we construct various measures, using 2–6 year citation counts. Two-year citation counts, thus, count publications that cite clinical trials published in the preceding two years. As our clinical trial sample begins in the year 2010, these measures are available only beginning in 2012. It is possible, and common, for clinical trials to cite other trials. We drop from our “citing” sample any records that we also flag as reporting the results of a trial.
4.1. Decomposing Types of Medical Research
@@
Figure 4.1 documents that the quantity of clinical trials—the black line—has remained essentially constant over the past ten years. In contrast, the quantity of other medical research—the colored lines—has increased substantially. Observe that we display the percent growth in the quantity of clinical trials in our sample published in each calendar year, alongside the growth in the set of papers that cite clinical trials, constructed in various ways to account for the truncation discussed above. Each series is normalized to begin at zero. This presentation masks substantial differences in the levels of these series. There are 10,903 clinical trials in our census in 2010, and 30,841 publications that cite clinical trials (using three-year citations) in 2013.
Figure 4.1. Growth in Clinical Research, Stability of Clinical Trials
@@
Notes: Figure 4.1 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year. Each series is reported in terms of the percent change relative to its initial value. The sample of published clinical trials is constructed with the conservative model. To address truncation, we report the number of publications that cite clinical trials published in the preceding $t$ years for each $t$ between 2 and 6.
These trends are robust to alternative forms of data construction. In Figure D.1, we construct these plots using our liberal and moderate sample definitions and find, qualitatively, no change. In Figure D.2, we explore a potential concern with a citation-based measure. We document in Figure 3.1 that the quantity of scientific publications is rising over time. It is reasonable to expect that this rise, coupled with the ease of citing papers enabled by the internet, coincides with an increase in the total number of citations in the aggregate and per paper. We confirm, empirically, that both hypotheses are borne out by our data. To account for the fact that the “cost” of citing a clinical trial may have decreased over time, thus generating the trend documented in Figure 4.1, we construct a version of this plot in which each citing paper is weighted by the total number of papers that it cites. That is, a citing paper that cites a trial and one other paper receives a weight of 1/2, and a citing paper that cites a trial and nine other papers receives a weight of 1/10. In doing so, we “penalize” papers that are especially generous in their tendency to cite other papers. Figure D.2 suggests, again, that this trend is robust to this alternate construction.
4.2. Clinical Trials as a Productivity Indicator
The trends in Figure 4.1, on their own, allow us to draw two conclusions. First, existing estimates that suggest sharply rising quantities of clinical trials (e.g., Bloom et al.,, 2020) are a consequence of technical challenges in the measurement of research. Second, existing estimates of rising quantities of medical and scientific research (e.g., Bloom et al.,, 2020; Chu and Evans,, 2021; Park et al.,, 2023) are not driven by shifts in the quantity of clinical trials.
Recall the very simple conceptual framework from Section 2. We assume, following e.g., Bloom et al., (2020), that the left-hand side of the “idea production function” (2.1) is constant in this setting. If clinical trials, as captured in our data, are a stable proxy for inputs $S_{t}$—in the sense that published trials capture the same amount of underlying research in each period—then an unchanging number of clinical trials implies that research productivity is constant. We return to the appropriate interpretation of rising trends in other medical research below. Here, we collect evidence that clinical trials are a stable proxy for inputs.
Figure 4.2 documents that—on two additional dimensions—clinical trial production has been stable over the past decade. In particular, Figure 4.2 displays the share of publications in our census with the following characteristics: any funding from a government agency (purple), any (three-year) citations from a leading journal in medicine (teal), or any citations from a leading journal conditional on having public funding (green).
Figure 4.2. Public Funding and Never-Cited Research